OpenStudy (anonymous):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Which of the following shows a proportion? A) 4/9 = 6/11 B) 3/4 = 8/6 C) 6/12 = 4/8 D) 9/15= 20/12

OpenStudy (anonymous):

those are fractions

jimthompson5910 (jim_thompson5910):

You have two basic options Option A Use a calculator: 4/9 = 0.4444 and 6/11 = 0.54545454 So 4/9 does NOT equal 6/11 Do the same with the others OR Option B 4/9 = 6/11 4*11 = 6*9 44 = 54 ... which is false, so 4/9 = 6/11 is also false

OpenStudy (anonymous):

6/12=4/8